OpenStudy (anonymous):

having trouble with this one find the coordinates of the midpoint of H[5 1/2, -4 1/4], X[3 3/4, -1 3/4]

OpenStudy (jdoe0001):

\(\bf \text{middle point of 2 points }\\ \left(\cfrac{x_2 + x_1}{2},\cfrac{y_2 + y_1}{2} \right)\)

OpenStudy (jdoe0001):

change the mixed numbers to single fractions and add them up

OpenStudy (anonymous):

(4.625,-3) is that correct

OpenStudy (austinl):

Typically, in mathematics you will want to leave your answers in exact form unless otherwise instructed to approximate into decimal form.

OpenStudy (jdoe0001):

well, using the calculator to add the fractions, I think is not what was expected

OpenStudy (austinl):

What?

OpenStudy (jdoe0001):

he didn't add the fractions, just used the calculator, thus the float

OpenStudy (anonymous):

I am trying to find the coordinates of the midpoint of HX

OpenStudy (austinl):

Ah, ok that makes more sense. Sorry, I guess I am just tired. \[H(5 \frac{1}{2}, -4 \frac{1}{4}), X(3 \frac{3}{4}, -1 \frac{3}{4})\] MIDPOINT: \[(\frac{x_1 + x_2}{2} , \frac{y_1 + y_2}{2}) = (\frac{5\frac{1}{2} + 3\frac{3}{4}}{2}, \frac{-4\frac{1}{4} +(-1\frac{3}{4})}{2})\]

OpenStudy (anonymous):

how do i simplify that

OpenStudy (austinl):

\[5\frac{1}{2} + 3\frac{3}{4} = \frac{11}{2} + \frac{15}{4}=\frac{22}{4}+\frac{15}{4}=\frac{37}{4}\rightarrow\frac{\frac{37}{4}}{2}=\frac{37}{8}\]

OpenStudy (anonymous):

the mid oint is a fraction its suposed to be an ordered pair

OpenStudy (austinl):

That was the first half of the ordered pair...

OpenStudy (anonymous):

oh ok second half -3?

OpenStudy (austinl):

Correct!